### v1

@@ 3 Method
In section 3.1, we define the STC task and then reiterate its primary challenges. Section 3.2 introduces Quartet Logic: A Four-step Reasoning Framework, a specialized CoT method designed for LLMs to address the STC task. Finally, we detail the CoT-Driven Multi-task learning method developed to enable smaller models to tackle the STC task effectively in section 3.3.
@@ 3.1 Task Description
Given a short text $X_{i}$ belonging to $D$ and a dataset consisting of N short texts, the STC task aims to classify $X_{i}$ as relevant label $l$ from a predefined set of labels $L$.
Semantic and syntactic challenges are inherent in STC tasks owing to their concise and informal nature. State-of-the-art methods based on GCN are constrained by their dependence on an external knowledge base. To address these limitations, we propose distinct frameworks for LLMs and smaller models that target the issues of limited context and reliance on external knowledge.
3.2 Quartet Logic: A Four-Step Reasoning Framework
To improve the performance of LLMs in STC tasks, this study introduces Quartet Logic: A Four-Step Reasoning (QLFR) framework, a CoT-prompt-based approach for multi-step reasoning. Rather than using conventional single-step prompting, which usually yields immediate label predictions without intermediary reasoning stages, we aimed to elicit inherent knowledge and abilities before determining the predicted label.
Our CoT method, called Syntactic and Semantic Enrichment CoT (SSE-CoT), provides a four-step reasoning approach, as shown in Figure 2. The first step involves identifying the key concepts within the input text $X_{i}$. Next, the method retrieves the relevant common sense knowledge $S$ embedded within the LLM. Subsequently, the text was revised to $R$ to enhance the clarity and context, enriching its interpretative depth. Finally, the revised text is used as a query to determine the predicted label, $\hat{y}$.
Figure 2: This diagram presents the Syntactic and Semantic Enrichment CoT (SSE-CoT), as applied to the short text ‘Del Potro says make French Open’. It begins by identifying key concepts, ‘Del Potro’ and the ‘French Open’, then combines them to contextualize ‘Del Potro’ as a tennis player and the ‘French Open’ as a major tournament. The third step refines this information for accuracy and integration. Finally, the process classifies the outcome under ‘sport’. The framework offers a novel solution that effectively addresses STC task challenges.
We conduct our four-step prompts as follows.
@@ 3.2 Quartet Logic: A Four-Step Reasoning Framework
In the initial step, a specified template is used to query the LLM and pose questions, utilizing its proficiency in identifying relevant concepts. $C^{1}_{1}$[Given the short text $X_{i}$], identify key concepts.
Here, $C^{1}_{1}$ comprises the context of the first step, and the following content constitutes the reasoning instruction $I^{1}_{1}$ for the first step. This stage was designed to focus the model on essential content in preparation for the next steps. The process can be formally expressed as:
@@
With the fundamental concepts $K^{1}$ identified in Step 1, this step involves retrieving the associated common-sense knowledge from the inherent knowledge base of the model using the following template: $C^{1}_{2}$[$C^{1}_{1}$,$K^{1}$], retrieve related common knowledge.
@@
Here, $f_{\mathrm{retrieve}}$ is a function enabling the model to recall pertinent information from an internal knowledge repository. Knowledge retrieval mitigates the semantic gap in STC tasks because of its brevity and facilitates the provision of contextually rich and comprehensive responses.
@@
Following the retrieval of pertinent common-sense knowledge, this step entails assimilating $S$ into a cohesive and polished short text. The context is formed by concatenating $C^{1}_{2}$ and $S$, where $I^{1}_{3}$ serves as the modification directive. $C^{1}_{3}$[$C^{1}_{2}$,$S$]. Refine and enhance the language to guarantee precision, fluidity, and legibility, whilst preserving the accuracy and wholeness of the integrated information.
Integration, represented by function $g$, is essential for converting raw data into an easily understandable format. It overcomes the syntactic constraints of short texts, producing a structured output $R$ that simplifies comprehension and subsequent categorization by LLM.
@@
Given the short text $R$. classify it into one of the categories. The categories are ‘health’, ‘sport’, ‘entertainment’, ‘business’, ‘sci_tech’, ‘U.S.’ and ‘world’.
After integrating common-sense knowledge and refining the short text $X_{i}$ as $R$, we prompted the LLM with instruction $I^{1}_{4}$ to generate the final predicted label.
$\hat{y_{i}}=\mathrm{argmax}p(y|X^{\prime},I_{4})$ (4)
@@
3.3 CoT-Driven Multi-task Learning Method
We propose a novel approach, the CoT-Driven Multi-Task learning (QLFR-CML) framework, specifically designed for smaller models to address STC tasks. The architecture of this framework is illustrated in Figure 3. Our framework consists of two stages, as detailed below.
Figure 3: Overview of the QLFR-CML method. In the first stage, the framework employs SSE-CoT and DA-CoT to prompt LLM with training data for rationale generation. In the second stage, the generated rationales guide the training of a smaller, specialized model. This stage involves multi-task fine-tuning that incorporates a supervised signal, which includes a label and two distinct rationales derived from SSE-CoT and DA-CoT reasoning.
In the initial stage, we employed two specialized CoT prompt processes to generate a rationale from an LLM. The Syntactic and Semantic Enrichment CoT (SSE-CoT) enhances the clarity and coherence of short texts by addressing their syntactic and semantic shortcomings. Simultaneously, Domain Augmentation CoT (DA-CoT) broadens the textual context by integrating domain-specific knowledge.
In the second stage, knowledge transfer occurs from the LLM to a smaller model through a multi-task learning strategy. The smaller model was trained to predict labels and accurately generate common-sense and domain-specific rationales. This integrated training approach improves the reasoning capabilities of the model and strengthens its ability to classify short texts with precision and depth.
@@ 3.3.1 Rationale generation
We introduce two distinct CoTs to generate rationales to enhance knowledge transfer in smaller models. First, SSE-CoT, introduced in Section LABEL:s3.2, addresses common-sense knowledge, encompassing broad, universally acknowledged information, such as fundamental life facts and recognized concepts. Second, Domain Augmentation CoT (DA-CoT) focuses on domain-specific knowledge, which includes detailed field-specific information, specialized jargon, and unique concepts. The DA-CoT uses a two-step reasoning process, as shown in Figure 4. We take the output $R$ from the third step of the SSE-CoT and the output $O$ from the second step of the DA-CoT as rationales.
@@
In the first step, we used a designated template to query the LLM, capitalizing on its capacity to pinpoint pertinent concepts. Unlike the first step of the SSE-CoT, the DA-CoT incorporates domain-specific cue words. For example, in the news domain illustrated in Figure. 4, it is essential to include critical entities, actions, and events. Details of the other domains are provided in Appendix A. $C^{2}_{1}$ comprises the context of the first step, and the following content constitutes the reasoning instruction $I^{2}_{1}$ for the first step.
$C^{2}_{1}$[Given the short text $X_{i}$], identify the key components, consider the main entities, actions, and events described.
@@
After establishing foundational concepts in the initial phase, this step prompts the LLM to apply domain-specific terminology and deeper analytical perspectives to its outputs. $C^{2}_{2}$[$C^{2}_{1}$,$K^{2}$]. Provide a summary of the identified components, including their interrelations and the overall significance within the context of the text.
@@ 3.3.3 Multi-task learning
The primary task is to predict the correct category label $\hat{y_{i}}$ from augmented input $X_{i}^{\prime}$. The secondary task involves processing the original text $X_{i}$ to output a rationale $\hat{r_{i}}$, with $r_{i}\in R$ serving as ground truth. Similarly, the tertiary task processes the original text $X_{i}$ to generate the domain-specific rationale $\hat{o_{i}}$, where $o_{i}\in O$ is the ground truth. This multi-task setup aims to predict the category label directly and generate rationales that provide interpretability and context to the model’s decisions, leveraging the rationales from the teacher model as supplementary guidance. Thus, the loss function encompasses the following terms for each task:
$L=L_{label}+\lambda_{1}L_{\text{SSE}}+\lambda_{2}L_{\text{DA}}$ (9)
@@
$L_{label}=\frac{1}{N}\sum_{i=1}^{N}\ell(f_{s}(X_{i}^{\prime}),\hat{y_{i}})$ (10)
$L_{SSE}=\frac{1}{N}\sum_{i=1}^{N}\ell(f_{s}(X_{i}),r_{i})$ (11)
$L_{DA}=\frac{1}{N}\sum_{i=1}^{N}\ell(f_{s}(X_{i}),o_{i})$ (12)
@@ 4.1 Datasets
To ensure a thorough and unbiased evaluation, we conducted extensive experiments on six widely recognized benchmark short-text datasets: MR, Snippets, Ohsumed, StackOverflow, TagMyNews, and AGNews. Following [35], we randomly sampled 40 labeled short texts from each class, where half formed the training set, and the other half formed the validation set. Table 1 provides detailed information regarding the datasets. We describe the datasets further below.
@@ 4.2 Experimental Setting
We selected LLaMA2-13B[40] as the foundational model for our QLFR method and chose the LLaMA2-13B and Flan-T5-Large[56] models to represent the LLM and smaller model in the QLFR-CML method, respectively. In the QLFR method, the LLaMA2-13B employs efficient parameter fine-tuning with Low-Rank Adaptation (LoRA)[57], which is conducted with a batch size of 10 across five epochs. LoRA maintains the weights of pretrained LMs while introducing trainable rank decomposition matrices into each transformer layer, making it feasible to fine-tune larger LMs with fewer computational resources [^7] In our experiment, trainable parameters only account for 0.24% of the entire LLaMA2-13B parameters. Conversely, in the QLFR-CML, Flan-T5-Large has fully fine-tuned parameters with batch sizes of 5 and 10 epochs. All experiments were conducted using three A100s and five V100s.
@@ 4.5 Benchmark Comparison
Table 2: Performance evaluation measured on short text datasets. This table presents the comparative performance of baseline models and ours, measured in terms of ACC (%) and F1 (%). The table highlights the highest scores in bold, with underscores indicating the highest scores achieved by prior methods. *** indicates that the result is reproduced by us.
Table 2 shows the performance comparison. As can be seen, our QLFR method surpasses other approaches on all six datasets. Similarly, the QLFR-CML method shows notable performance across five datasets. Specifically, on the TagMyNews dataset, QLFR achieves an increase of 12.60% in ACC and 12.05% in F1 score compared to the previously optimal ChatGLM model. QLFR-CML records an 8.84% improvement in ACC and a 7.64% increase in F1 score. These results support our hypothesis that LLMs can effectively utilize their inherent knowledge and abilities to address traditional NLP tasks, particularly the STC task discussed in this paper, thereby validating the effectiveness of our proposed methods.
A comparative analysis revealed that QLFR outperformed QLFR-CML, especially on the MR dataset, where QLFR’s accuracy of QLFR exceeded that of QLFR-CML by 5.6%. This indicates that models with larger parameters have greater intrinsic knowledge and enhanced capabilities to solve the STC task. Although the CML did not outperform ChatGLM on the StackOverflow dataset, it significantly exceeded the top-performing model in the GCN-based group. Moreover, it achieved a 7.3% higher ACC and a 7.77% greater F1 score than SHINE.

### v2

@@ 3 Method
In section 3.1, we define the STC tasks and then reiterate its primary challenges. Section 3.2 introduces SSE-CoT, a specialized CoT method designed for LLMs to better address the STC tasks. Finally, we detail the CoT-Driven Multi-Task learning method developed to enable smaller models to tackle the STC tasks effectively in section 3.3.
@@ 3.1 Task Description
Given a dataset $D$ consisting of N short texts, the STC tasks aim to classify text $x_{i}\in D$ as a relevant label $l$ from a predefined set of labels $L$. Semantic and syntactic challenges are inherent in STC tasks owing to their concise and informal nature. State-of-the-art methods based on GCN are constrained by their dependence on an external knowledge base. To address these limitations, we propose distinct methods for LLMs and smaller models that target the issues of limited context and dependency on external knowledge.
3.2 Semantic and Syntactic Enrichment CoT in LLMs
To enhance the performance of LLMs in handling STC tasks, this study introduces the Semantic and Syntactic Enrichment CoT (SSE-CoT). Unlike traditional methods that employ single-step prompts, SSE-CoT employs a multi-step reasoning process specifically designed to enhance both the semantic and syntactic understanding of short texts within LLMs.
  • Semantic elements are concerned with the meanings that words, phrases, and sentences convey through the short texts. They involve understanding the implications, nuances, and contextual uses of language within a narrative.
  • Syntactic elements refer to the arrangement of words and phrases to create well-formed sentences according to the rules of grammar. This includes the structure of sentences, the correct use of grammatical rules, and the logical relationships between different parts of the text.
Our SSE-CoT method comprises four distinct steps, as shown in Fig 2. First, it identifies key concepts and terms to establish a foundational understanding. The second step, semantic enrichment, deepens the model’s comprehension of these concepts and their relationships. This is followed by syntactic enrichment, which refines the grammar and structure of the text. The final step integrates these enhancements, enabling the model to make predictions. Specifically, we conduct our four-step prompts as follows.
Figure 2: This diagram presents the Semantic and Syntactic Enrichment CoT (SSE-CoT), as applied to the short text ‘Del Potro says make French Open’. It begins by identifying key concepts, ‘Del Potro’ and the ‘French Open’, then combines them to contextualize ‘Del Potro’ as a tennis player and the ‘French Open’ as a major tournament. The third step refines this information for accuracy and integration. Finally, the process classifies the outcome under ‘sport’. The framework offers a novel solution that effectively addresses STC tasks challenges.
@@ 3.2 Semantic and Syntactic Enrichment CoT in LLMs
We first ask LLM to identify relevant concepts using a specified template: {mdframed}[ outerlinewidth=0.5pt, roundcorner=5pt, innertopmargin=5pt, innerbottommargin=5pt, innerrightmargin=5pt, innerleftmargin=5pt, backgroundcolor=gray!7, linecolor=black, align=center, userdefinedwidth=0.5] $C^{1}_{1}$[Given the short text $x_{i}$], identify key concepts.
Here, $C^{1}_{1}$ is the context of the first step, and the following content is the instruction $I^{1}_{1}$. This stage was designed to focus the model on essential content in preparation for the next steps. The process can be formally expressed as:
@@
With the fundamental concepts $K^{1}$ identified in Step 1, this step involves retrieving the associated common-sense knowledge from the inherent knowledge base of the LLM using the following template: {mdframed}[ outerlinewidth=0.5pt, roundcorner=5pt, innertopmargin=5pt, innerbottommargin=5pt, innerrightmargin=5pt, innerleftmargin=5pt, backgroundcolor=gray!7, linecolor=black, align=center, userdefinedwidth=0.5] $C^{1}_{2}$[$C^{1}_{1}$,$K^{1}$], retrieve related common knowledge.
@@
here, $f_{\mathrm{retrieve}}$ is a function enabling the model to recall pertinent information from an internal knowledge repository. Knowledge retrieval mitigates the semantic gap in STC tasks because of its brevity, and facilitates the provision of contextually rich and comprehensive responses.
@@
Following the retrieval of pertinent common-sense knowledge, this step entails assimilating $S$ into a cohesive and polished short text. The context is formed by concatenating $C^{1}_{2}$ and $S$, where $I^{1}_{3}$ serves as the modification directive. {mdframed}[ outerlinewidth=0.5pt, roundcorner=5pt, innertopmargin=5pt, innerbottommargin=5pt, innerrightmargin=5pt, innerleftmargin=5pt, backgroundcolor=gray!7, linecolor=black, align=center, userdefinedwidth=0.5] $C^{1}_{3}$[$C^{1}_{2}$,$S$]. Refine and enhance the language to guarantee precision, fluidity, and legibility, whilst preserving the accuracy and wholeness of the integrated information.
Integration, represented by function $g$, is essential for converting raw data into an easily understandable format. It overcomes the syntactic constraints of short texts, producing a structured output $R$ that simplifies comprehension and subsequent categorization by LLM. The process can be formally expressed as:
@@
After integrating common-sense knowledge and refining the short text $x_{i}$ as $R$, we prompted the LLM with instruction $I^{1}_{4}$ to generate the final predicted label. {mdframed}[ outerlinewidth=0.5pt, roundcorner=5pt, innertopmargin=5pt, innerbottommargin=5pt, innerrightmargin=5pt, innerleftmargin=5pt, backgroundcolor=gray!7, linecolor=black, align=center, userdefinedwidth=0.5] Given the short text $R$. classify it into one of the categories. The categories are ‘health’, ‘sport’, ‘entertainment’, ‘business’, ‘sci_tech’, ‘U.S.’ and ‘world’.
The process can be formally expressed as:
$\hat{y_{i}}=\mathrm{argmax}p(y|R,I_{4})$ (4)
@@
3.3 CoT-Driven Multi-Task learning for Smaller Models
We propose the CoT-Driven Multi-Task learning (CDMT) method for STC tasks using smaller models. The architecture of this framework is depicted in Fig 3. Our framework comprises two stages, described below.
Figure 3: Overview of the CDMT method. In the first stage, the framework employs SSE-CoT and DA-CoT to prompt LLM with training data for rationale generation. In the second stage, the generated rationales guide the training of a smaller, specialized model. This stage involves multi-task fine-tuning that incorporates a supervised signal, which includes a label and two distinct rationales derived from SSE-CoT and DA-CoT reasoning.
In the initial stage, we employed two specialized CoT prompt processes to generate rationales from LLM. The SSE-CoT enhances the clarity and coherence of short texts by addressing their inherent limitations. We define a rationale as a chain of reasoning processes generated by LLMs in this paper. Concurrently, Domain Augmentation CoT (DA-CoT) enriches the textual context by incorporating domain-specific knowledge. In the second stage, knowledge transfer occurs from the LLM to a smaller model through a multi-task learning strategy. The smaller model was trained to predict labels and generate common-sense and domain-specific rationales. This integrated training approach improves the reasoning capabilities of the model and strengthens its ability to classify short texts with precision and depth.
@@ 3.3.1 Rationale generation
In our study, we introduce two distinct CoTs to generate rationales. First, SSE-CoT introduced in Section 3.2,SSE-CoT is specifically designed for STC tasks to address the unique characteristics of short texts directly. Second, DA-CoT aims to enhance the performance of smaller models by transferring more domain knowledge. This approach follows a two-step reasoning process, as shown in Figure 4. Specifically, we conduct DA-CoT prompts as follows.
@@
In the first step, we used a designated template to query the LLM, capitalizing on its capacity to pinpoint pertinent concepts. Unlike the first step of the SSE-CoT, the DA-CoT incorporates domain-specific cue words. For example, in the news domain illustrated in Fig 4, it is essential to include critical entities, actions, and events. Details of the other domains are provided in Appendix A. $C^{2}_{1}$ comprises the context of the first step, and the following content constitutes the reasoning instruction $I^{2}_{1}$ for the first step.
{mdframed}
[ outerlinewidth=0.5pt, roundcorner=5pt, innertopmargin=5pt, innerbottommargin=5pt, innerrightmargin=5pt, innerleftmargin=5pt, backgroundcolor=gray!7, linecolor=black, align=center, userdefinedwidth=0.5] $C^{2}_{1}$[Given the short text $x_{i}$], identify the key components, consider the main entities, actions, and events described.
@@
After establishing foundational concepts in the initial phase, this step prompts the LLM to apply domain-specific terminology and deeper analytical perspectives to its outputs. {mdframed}[ outerlinewidth=0.5pt, roundcorner=5pt, innertopmargin=5pt, innerbottommargin=5pt, innerrightmargin=5pt, innerleftmargin=5pt, backgroundcolor=gray!7, linecolor=black, align=center, userdefinedwidth=0.5] $C^{2}_{2}$[$C^{2}_{1}$,$K^{2}$]. Provide a summary of the identified components, including their interrelations and the overall significance within the context of the text.
@@
The rationales generated by SSE-CoT directly influence the classification outcomes, whereas those from DA-CoT primarily enhance performance implicitly. Thus, it is necessary to filter and verify the rationales produced by SSE-CoT. Following the CROP method [53], We input $x_{i}$ with SSE-CoT template into the LLM to obtain the intermediate explanation $R$ and the predicted label $\hat{y_{i}}$. We accept $r_{i}=R$ only if $y_{i}=\hat{y_{i}}$. If they do not match, we concatenate $x_{i}$ with its true label $y_{i}$ to generate $R$ without applying any filter.
@@ 3.3.3 Multi-task learning
The primary task is to predict the correct category label $\hat{y_{i}}$ from augmented input $x_{i}^{\prime}$. The secondary task involves processing the original text $x_{i}$ to output a rationale $\hat{r_{i}}$, with $r_{i}$ defined in Section 3.3.1 serving as ground truth. Similarly, the tertiary task processes the original text $x_{i}$ to generate the domain-specific rationale $\hat{o_{i}}$, where $o_{i}\in O$ is the ground truth. This multi-task setup aims to predict the category label directly and generate rationales that provide interpretability and context to the model’s decisions, leveraging the rationales from the teacher model as supplementary guidance. Thus, the loss function encompasses the following terms for each task:
$L=L_{\text{label}}+\lambda_{1}L_{\text{SSE}}+\lambda_{2}L_{\text{DA}}$ (9)
@@
$L_{\text{label}}=\frac{1}{N}\sum_{i=1}^{N}\ell(f_{s}(x_{i}^{\prime}),\hat{y_{i}})$ (10)
$L_{\text{SSE}}=\frac{1}{N}\sum_{i=1}^{N}\ell(f_{s}(x_{i}),r_{i})$ (11)
$L_{\text{DA}}=\frac{1}{N}\sum_{i=1}^{N}\ell(f_{s}(x_{i}),o_{i})$ (12)
@@
  • The primary loss function, $L_{\text{label}}$, ensures that the model correctly classifies the short texts into their respective categories.
  • The second, $L_{\text{SSE}}$, relates to the SSE-CoT rationale generation, ensuring that the model not only performs well on classification but also aligns its reasoning with how the SSE-CoT enhances short text understanding.
  • The third, $L_{\text{DA}}$, associated with the DA-CoT, ensures that domain-specific knowledge is incorporated effectively.
Optimizing short text classification, text refinement, and related knowledge prediction within a multi-task framework effectively utilizes the induced knowledge of LLMs. By incorporating moderate inductive biases into the parameter space, this approach enhances the generalization performance and robustness of smaller models.
Algorithm 1 CDMT Framework
1:Input: Samples $X=\{x_{1},x_{2},\dots,x_{n}\}$ and labels $Y=\{y_{1},y_{2},\dots,y_{n}\}$triangleright First Stage: Rationale Generation with LLMs
2:for each $(x_{i},y_{i})$ in $(X,Y)$ do
3:     Generate $R,\hat{y}_{i}$ using SSE-CoT with $x_{i}$
4:     if $\hat{y}_{i}=y_{i}$ then
5:         Set $r_{i}=R$
6:     else
7:         Concatenate $x_{i}$ with true label $y_{i}$
8:         Regenerate $R,\hat{y}_{i}$ using SSE-CoT with $x_{i},y_{i}$
9:         Set $r_{i}=R$
10:     end if
11:     Generate $O$ using DA-CoT with $x_{i}$
12:     Set $o_{i}=O$
13:end $\triangleright$ Second Stage: Fine-tuning Smaller Model
14:Initialize smaller model
15:for each $(x_{i},r_{i},y_{i},o_{i})$ do
16:     Prepare $x^{\prime}_{i}=l_{1}\oplus l_{2}\oplus\ldots\oplus l_{m}\oplus x_{i}$
17:     Train smaller model using multi-task learning:
18:      - Compute loss $L_{\text{label}}$ based on $\ell(f_{s}(x^{\prime}_{i}),y_{i})$
19:      - Compute loss $L_{\text{SSE}}$ based on $\ell(f_{s}(x_{i}),r_{i})$
20:      - Compute loss $L_{\text{DA}}$ based on $\ell(f_{s}(x_{i}),O_{i})$
21:      - Update model by minimizing $L=L_{\text{label}}+\lambda_{1}L_{\text{SSE}}+\lambda_{2}L_{\text{DA}}$
22:end for
23:Output: Fine-tuned smaller model
@@ 4.1 Datasets
To ensure a thorough and unbiased evaluation, we conducted extensive experiments on six widely recognized benchmark short-text datasets: MR, Snippets, Ohsumed, StackOverflow, TagMyNews, and AGNews. Following [35], we randomly sampled 40 labeled short texts from each class, where half formed the training set, and the other half formed the validation set. Table 1 provides detailed information regarding the datasets. We describe the datasets in more detail further below.
@@ 4.2 Experimental Setting
We selected LLaMA2-13B[40] as the foundational model for our SSE-CoT method and chose the LLaMA2-13B and Flan-T5-Large[58] models to represent the LLM and smaller model in the CDMT method, respectively. In the SSE-CoT method, the LLaMA2-13B employs efficient parameter fine-tuning with Low-Rank Adaptation (LoRA)[59], which is conducted with a batch size of 10 across five epochs. LoRA maintains the weights of pretrained LMs while introducing trainable rank decomposition matrices into each transformer layer, making it feasible to fine-tune larger LMs with fewer computational resources [^8] In our experiment, trainable parameters only account for 0.24% of the entire LLaMA2-13B parameters. Conversely, in the CDMT, Flan-T5-Large has fully fine-tuned parameters with batch sizes of 5 and 10 epochs. All experiments were conducted using three A100s and five V100s.
@@ 4.5 Benchmark Comparison
Table 2: Performance evaluation measured on short text datasets. This table presents the comparative performance of baseline models and ours, measured in terms of ACC (%) and F1 (%). The table highlights the highest scores in bold, with underscores indicating the highest scores achieved by prior methods. * indicates that the result is reproduced by us.
Table 2 shows the performance comparison. As can be seen, our SSE-CoT method surpasses other approaches on all six datasets. Similarly, the CDMT method shows notable performance across five datasets. Specifically, on the TagMyNews dataset, SSE-CoT achieves an increase of 12.60% in ACC and 12.05% in F1 score compared to the previously optimal ChatGLM model. CDMT records an 8.84% improvement in ACC and a 7.64% increase in F1 score. These results support our hypothesis that LLMs can effectively utilize their inherent knowledge and abilities to address traditional NLP tasks, particularly the STC tasks discussed in this paper, thereby validating the effectiveness of our proposed methods.
A comparative analysis revealed that SSE-CoT outperformed CDMT, especially on the MR dataset, where SSE-CoT’s accuracy of SSE-CoT exceeded that of CDMT by 5.6%. This indicates that models with larger parameters have greater intrinsic knowledge and enhanced capabilities to solve the STC tasks. Although the CML did not outperform ChatGLM on the StackOverflow dataset, it significantly exceeded the top-performing model in the GCN-based group. Moreover, it achieved a 7.3% higher ACC and a 7.77% greater F1 score than SHINE.